# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 5.A. Trees associated with an automorphism of a one-ended relatively hyperbolic group
Note that, since $v$ is rigid or QH with fibre, the preimages of $e_{1}$ and $e_{2}$ in $T_{0}^{\phi}$ are unique edges since $T^{\phi}$ is obtained from $T_{0}^{\phi}$ by collapsing edges with elementary endpoints. Thus, it suffices to prove the result for $T_{0}^{\phi}$.
Let $v\in VT_{0}^{\phi}$ be either rigid or QH with fibre and let $e_{1},e_{2}\in ET_{0}^{\phi}$ be two distinct edges adjacent to $v$. Suppose first that both $e_{1}$ and $e_{2}$ are edges in either $T^{\mathrm{can}}$ or $T_{w}$ for some $w\in VT^{\mathrm{can}}$ which is QH with fibre. Then $G_{e_{1}}\cap G_{e_{2}}$ is finite and $\langle G_{e_{1}},G_{e_{2}}\rangle$ is not elementary by 4.2 $(8)$.
Thus, we may suppose, up to reordering, that $e_{1}$ is an edge coming from $T^{\mathrm{can}}$ and $e_{2}$ is an edge coming from some $T_{w}$ with $w\in VT^{\mathrm{can}}$ QH with fibre. In particular, since $w$ is QH with fibre, the group $G_{e_{1}}$ is virtually cyclic.
Note that $e_{1}$ is attached to the centre of the fixed point set of $G_{e_{1}}$ in $T_{w}$. Moreover, by 4.2 $(8)$, this centre is not elementary if and only if the fixed point set is reduced to a point. Thus, the fixed point set of $G_{e_{1}}$ in $T_{w}$ is reduced to $v$ (seen as a point in $T_{w}$).
Suppose towards a contradiction that $G_{e_{1}}\cap G_{e_{2}}$ is infinite and let $v^{\prime}$ be the endpoint of $e_{2}$ distinct from $v$. As explained above, the point $v^{\prime}$ is not fixed by $G_{e_{1}}$. Let $g\in G_{e_{1}}$ be such that $gv^{\prime}\neq v^{\prime}$. Then one of the endpoints of $ge_{2}$ is $v$ since $G_{e_{1}}$ fixes $v$. Moreover, since $G_{e_{1}}$ is virtually cyclic, the intersection $G_{e_{2}}\cap gG_{e_{2}}g^{-1}$ is infinite (it contains the normal core in $G_{e_{1}}$ of the intersection $G_{e_{1}}\cap G_{e_{2}}$). Therefore, $e_{2}$ and $ge_{2}$ are two distinct edges of $T_{w}$ adjacent to the nonelementary vertex $v$ such that $G_{e_{2}}\cap gG_{e_{2}}g^{-1}$ is infinite. This contradicts 4.2 $(8)$. Thus, the intersection $G_{e_{1}}\cap G_{e_{2}}$ is finite.
We now prove that $\langle G_{e_{1}},G_{e_{2}}\rangle$ is not elementary. Since $w$ is a QH with fibre vertex of $T^{\mathrm{can}}$, the group $G_{w}$ is Gromov hyperbolic. Therefore, the only elementary subgroups of $G$ contained in $G_{w}$ are virtually cyclic. Since $G_{e_{1}}\cap G_{e_{2}}$ is finite, the group $\langle G_{e_{1}},G_{e_{2}}\rangle$ is not virtually cyclic, hence is not elementary.
The fact that $T^{\phi}$ is $2$-acylindrical follows from the fact that any edge of $T^{\phi}$ has an endpoint which is nonelementary. ∎
@@
Let $v\in T^{\phi}$ be rigid. The map $\langle\phi\rangle^{0}\to\operatorname{\mathrm{Out}}(G_{v})$ has finite image.
@@
Since $v$ is rigid, the group $G_{v}$ is the stabiliser of a rigid vertex of either $T^{\mathrm{can}}$ or some $T_{w}$ where $w\in VT^{\mathrm{can}}$ is QH with finite fibre.
If $G_{v}$ is the stabiliser of a rigid vertex of $T^{\mathrm{can}}$, then, by Theorem 4.2 $(10)$, then map $\langle\phi\rangle^{0}\to\operatorname{\mathrm{Out}}(G_{v})$ has finite image.
Suppose now that $G_{v}$ is the stabiliser of a rigid vertex of some $T_{w}$ where $w\in VT^{\mathrm{can}}$ is QH with finite fibre. By Lemma 5.2, the set $\mathrm{Per}_{\mathrm{NP}}([\phi|_{G_{w}}])\cup\{[G_{e_{1}}],\ldots,[G_{e_{k}}]\}$ is a finite set of conjugacy classes of finitely generated subgroups of $T_{w}$. Thus, Theorem 4.2 $(10)$ applies (to $T_{w}$) and the image of $\langle\phi\rangle^{0}\to\operatorname{\mathrm{Out}}(G_{v})$ is finite. ∎
@@
Let $v\in T^{\phi}$ and let $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$.
@@
Suppose first that $v$ is QH with fibre. By construction of $T^{\phi}$, the group $G_{v}$ is the stabiliser of a QH with fibre vertex of some JSJ tree $T_{w}$, where $w$ is a QH with fibre vertex of $T^{\mathrm{can}}$. But the intersection of $G_{v}$ with every subgroup $H^{\prime}$ such that $[H^{\prime}]\in\mathrm{Per}_{\mathrm{NP}}([\phi|_{G_{v}}])\cup\{[G_{e_{1}}],\ldots,[G_{e_{k}}]\}$ is elementary by the definition of QH with fibre vertices. Thus, the intersection of $H$ with $G_{v}$ is elementary. This proves Assertion $(1)$.
Suppose now that $v$ is rigid.
Let $\Psi\in\phi$ be such that $\mathrm{Per}(\Psi)=H$. Suppose that the intersection $H\cap G_{v}$ is nonelementary. Then $\Psi$ preserves $G_{v}$ since $v$ is the unique vertex in $T^{\phi}$ fixed by $H\cap G_{v}$. By Lemma 5.4, there exists $k\in\mathbb{N}^{*}$ such that $\Psi^{k}$ acts as a global conjugation on $G_{v}$ by an element $g\in G_{v}$. Taking a larger $k$ if necessary, we may also assume that $\Psi^{k}$ also act trivially on a nonelementary subgroup $H^{\prime}\subseteq H\cap G_{v}$.
Note that $\Theta=\mathrm{ad}_{g^{-1}}\circ\Psi^{k}$ acts trivially on $G_{v}$. Thus, $\Psi^{k}$ and $\Theta$ acts trivially on the same nonelementary subgroup $H^{\prime}\subseteq G_{v}$. By Lemma 5.1, there exists $N\in\mathbb{N}$ such that $\Psi^{kN}=\Theta^{N}$. In particular, $\Psi$ has a power which fixes $G_{v}$ elementwise. Thus, we have $G_{v}\subseteq H$. ∎
@@
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$ and let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. The rest of the section is dedicated to the proof of some properties of the set $\mathrm{Per}_{\mathrm{NP}}(\phi)$ and of the action of $G_{\Phi}$ on $T^{\phi}$. To this end we prove that $\mathrm{Per}_{\mathrm{NP}}(\phi)$ is finite (see Lemma 5.11). We need the following lemmas regarding the intersection of characteristic sets of isometries in $T^{\phi}$.
@@
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$, let $v\in VT^{\phi}$ be non-elementary and let $e_{1}$ and $e_{2}$ be distinct edges adjacent to $v$. Let $\phi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. If there is a representative $\Phi\in\phi$ such that $\Phi(G_{e_{i}})=G_{e_{i}}$ for each $i=1,2$, then $v$ is rigid and $G_{v}$ is fixed elementwise by some power of $\Phi$.
@@
We first prove that $v$ is rigid. Indeed, suppose towards a contradiction that $v$ is QH with finite fibre. Then $G_{e_{1}}$ and $G_{e_{2}}$ are virtually cyclic. Thus, for every $i\in\{1,2\}$, the automorphism $\Phi$ has a power $\Phi^{k}$ fixing an infinite order element $g_{i}\in G_{e_{i}}$. By Lemma 5.3 and as the groups $G_{e_{1}}$ and $G_{e_{2}}$ are virtually cyclic, the group $\langle g_{1},g_{2}\rangle$ is a non-elementary subgroup. As $\langle g_{1},g_{2}\rangle\subseteq\mathrm{Per}(\Phi)\cap G_{v}$, this contradicts Lemma 5.5. Thus, the vertex $v$ is rigid.
By Lemma 5.4, after taking a power $\Phi^{\ell}$, it acts on $G_{v}$ as global conjugation by an element $g\in G_{v}$.
We claim that, after taking a further power of $\Phi$, the element $g$ is trivial. Indeed, note that, by Lemma 5.3 the stabiliser of an edge adjacent to $v$ is almost malnormal in $G_{v}$: for every edge $e^{\prime}$ of $T^{\phi}$ adjacent to $v$ and every $g^{\prime}\in G_{v}$, if $g^{\prime}G_{e}g^{\prime-1}\cap G_{e}$ is infinite then $g^{\prime}\in G_{e}$. Moreover, if $e$ and $e^{\prime}$ are two distinct edges adjacent to $v$, then $G_{e}\cap G_{e^{\prime}}$ is finite. Since $\Phi$ preserves $G_{e_{1}}$ and $G_{e_{2}}$, the power $\Phi^{\ell}$ must act by conjugating by an element of the finite intersection $G_{e_{1}}\cap G_{e_{2}}$. This becomes trivial on taking a further power of $\Phi$, which concludes the proof of the claim.
@@
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$ and let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. Let $[H]=[\mathrm{Per}(\Phi)]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ and let $T_{H}^{\phi}$ be the minimal $H$-invariant subtree of $T^{\phi}$.
@@
Since $T_{H}^{\phi}$ contains an edge, it follows that $H$ is not elliptic in $T^{\phi}$. Thus, $T^{\phi}_{H}$ is the union of the axes of elements of $H$. Then, $e$ is contained in the axis of an element $g$ of $H$. Recalling the bipartite structure of $T^{\phi}$, let $v$ be the rigid or QH vertex adjacent to $e$. Then there exists an edge $e^{\prime}\neq e$ adjacent to $v$ and contained in the axis of $g$. Let $\Phi\in\phi$ be such that $\mathrm{Per}(\Phi)=H$ and let $N\in\mathbb{N}^{*}$ be such that $\Phi^{N}(g)=g$. Since $\Phi^{N}$ is elliptic in $T^{\phi}$ by Remark 5.6, it fixes pointwise the axis of $g$. In particular, it fixes $e$ and $e^{\prime}$. By Lemma 5.7, we see that $v$ is rigid and that $G_{v}$ is fixed elementwise by a power of $\Phi$. ∎
@@
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$ and let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. Let $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ and let $T_{H}^{\phi}$ be the minimal $H$-invariant subtree of $T^{\phi}$.
@@
Suppose first that $T_{H}^{\phi}$ is reduced to a point $v$. Then $H\subseteq G_{v}$ and $v$ is not QH with fibre by Lemma 5.5. Suppose now that $T_{H}^{\phi}$ is not reduced to a point. Then any vertex $v$ of $T_{H}^{\phi}$ is adjacent to an edge and the result follows from Corollary 5.8. ∎
@@
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$ and let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. Let $n\in\mathbb{N}^{*}$ and let $\Psi,\Theta\in\phi^{n}$ be such that $F_{\Psi}$ and $F_{\Theta}$ are elliptic isometries of $T^{\phi}$. Suppose that there exist $g,h\in G$ loxodromic in $T^{\phi}$, such that $\Psi(g)=g$, $\Theta(h)=h$ and $\mathrm{Ax}(g)\cap\mathrm{Ax}(h)$ contains an edge $e$.
@@
Therefore, we can apply Lemma 5.7: there exists $N\in\mathbb{N}^{*}$ such that both $\Psi^{N}$ and $\Theta^{N}$ act as the identity on the nonelementary subgroup $G_{v}$. By Lemma 5.1, up to taking powers of $\Psi$ and $\Theta$, we have $\Psi^{N}=\Theta^{N}$. ∎
@@
Let $G$ be a one-ended relatively hyperbolic group and let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G)$. The set $\bigcup_{n\in\mathbb{N}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ is finite. Consequently, there exists $N\in\mathbb{N}^{*}$ such that any subgroup whose conjugacy class is in $\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ belongs to a subgroup whose conjugacy class is in $\mathrm{Per}_{\mathrm{NP}}(\phi^{N})$.
@@
Counting elliptic subgroups: Let $[H]\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$. Suppose first that every element of $H$ is elliptic in $T^{\phi}$. We claim that $H$ is elliptic in $T^{\phi}$. Indeed, otherwise, by acylindricity of the action, there would exist two distinct $g,h\in H$ with $\operatorname{\mathrm{Fix}}(g)\cap\operatorname{\mathrm{Fix}}(h)=\varnothing$. Therefore, the element $gh\in H$ would be loxodromic by [CM87, Proposition 1.5], a contradiction. Therefore the group $H$ is elliptic in $T^{\phi}$.
Since $[H]\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$, the group $H$ is not contained in the stabiliser of a vertex which is elementary. By Corollary 5.8, it is also not contained in the stabiliser of a QH with finite fibre vertex. Therefore, the group $H$ is contained in the stabiliser of a vertex $v$ which is rigid. By Lemma 5.5, we have in fact $H=G_{v}$.

### 02 Mar 2026 v2

@@ 5.A. Trees associated with an automorphism of a one-ended relatively hyperbolic group
This is Theorem 4.2 $(2),(8)$. ∎
Recall that we denote by $\langle\phi\rangle^{0}$ the finite index subgroup of $\langle\phi\rangle$ which acts trivially on the quotient graph $G\backslash T^{\phi}$. For every $v\in VT^{\phi}$, we then have a well-defined homomorphism $\langle\phi\rangle^{0}\to\operatorname{\mathrm{Out}}(G_{v})$.
@@
Let $v\in VT^{\phi}$ be rigid. The map $\langle\phi\rangle^{0}\to\operatorname{\mathrm{Out}}(G_{v})$ has finite image.
@@
By Lemma 5.2 $(1)$, ${\mathcal{H}}_{QH}$ is a finite set of finitely generated groups, each periodic for some representative of a power of $\phi$, and so there exists $k>1$ such that $\phi^{k}\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)}\cup{\mathcal{H}}_{QH}^{(t)})$. The result now follows from Theorem 4.2 $(10)$. ∎
Recall that $T^{\mathrm{can}}$ and $T^{\phi}$ are compatible. Let $T^{\mathrm{ref}}_{0}$ be their minimal refinement. Since both $T^{\mathrm{can}}$ and $T^{\phi}$ are $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees, so is $T^{\mathrm{ref}}_{0}$.
@@
The tree $T^{\mathrm{ref}}_{0}$ is a JSJ tree over elementary subgroups relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH}$. In particular, $T^{\mathrm{ref}}_{0}$ and $T^{\phi}$ have the same set of elliptic subgroups.
Proof.
We check [30, Definition 2.11]. We first prove that $T^{\mathrm{ref}}_{0}$ is universally elliptic in $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees, that is every edge group of $T^{\mathrm{ref}}_{0}$ is elliptic in every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees. Let $G_{e}$ be an edge group of $T^{\mathrm{ref}}_{0}$. By minimality of $T^{\mathrm{ref}}_{0}$, the group $G_{e}$ is an edge group of either $T^{\mathrm{can}}$ or $T^{\phi}$. As $T^{\mathrm{can}}$ is a JSJ $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0})$-tree, it is universally elliptic in $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0})$-trees, hence it is universally elliptic in $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees. Similarly, $T^{\phi}$ is universally elliptic in $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees. Thus, $G_{e}$ is elliptic in every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees and $T^{\mathrm{ref}}_{0}$ is universally elliptic in $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees.
We now prove that $T^{\mathrm{ref}}_{0}$ dominates every universally elliptic $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees, that is, every vertex group of $T^{\mathrm{ref}}_{0}$ is elliptic in every universally elliptic $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-tree. Since $T^{\phi}$ is a JSJ $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-tree, it dominates every universally elliptic $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees. Since $T^{\mathrm{ref}}_{0}$ collapses onto $T^{\phi}$, every vertex group of $T^{\mathrm{ref}}_{0}$ is elliptic in $T^{\phi}$ hence is elliptic in every universally elliptic $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-tree. This shows that $T^{\mathrm{ref}}_{0}$ dominates every universally elliptic $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-trees. Together with the above paragraph, we obtain that $T^{\mathrm{ref}}_{0}$ is a JSJ tree over elementary subgroups relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH}$. ∎
Lemma 5.6.
Let $v\in T^{\phi}$ and let $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$.
@@
Suppose first that $v$ is QH with fibre. Suppose towards a contradiction that $H\cap G_{v}$ is nonelementary. By Lemma 5.5, the group $G_{v}$ is an elliptic group of the refined tree $T^{\mathrm{ref}}_{0}$. We still denote by $v$ the vertex in $T^{\mathrm{ref}}_{0}$ fixed by $G_{v}$. Let $p\colon T^{\mathrm{ref}}_{0}\to T^{\mathrm{can}}$ be the natural projection. Since $G_{v}$ is a flexible vertex, the image $p(v)$ is not a rigid vertex of $T^{\mathrm{can}}$ (recall that every rigid vertex stabiliser of $T^{\mathrm{can}}$ is also elliptic in every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH})$-tree). Hence $p(v)$ is a flexible vertex of $T^{\mathrm{can}}$. Therefore, if $H\cap G_{v}$ is nonelementary, then $H\cap G_{v}\in{\mathcal{H}}_{QH}$. This contradicts Theorem 4.2 $(9)$. Hence $H\cap G_{v}$ is elementary.
Suppose now that $v$ is rigid. Let $\Psi\in\phi$ be such that $\mathrm{Per}(\Psi)=H$. Suppose that the intersection $H\cap G_{v}$ is nonelementary. Then $\Psi$ preserves $G_{v}$ since $v$ is the unique vertex in $T^{\phi}$ fixed by $H\cap G_{v}$. By Lemma˜5.4, there exists $k\in\mathbb{N}^{*}$ such that $\Psi^{k}$ acts as a global conjugation on $G_{v}$ by an element $g\in G_{v}$. Taking a larger $k$ if necessary, we may also assume that $\Psi^{k}$ acts trivially on a nonelementary subgroup $H^{\prime}\subseteq H\cap G_{v}$.
Note that $\Theta=\mathrm{ad}_{g^{-1}}\circ\Psi^{k}$ acts trivially on $G_{v}$. Thus, $\Psi^{k}$ and $\Theta$ acts trivially on the same nonelementary subgroup $H^{\prime}\subseteq G_{v}$. By Lemma˜5.1, there exists $N\in\mathbb{N}$ such that $\Psi^{kN}=\Theta^{N}$. In particular, $\Psi$ has a power which fixes $G_{v}$ elementwise. Thus, we have $G_{v}\subseteq H$. ∎
@@
The rest of the section is dedicated to the proof of some properties of the set $\mathrm{Per}_{\mathrm{NP}}(\phi)$ and of the action of $G_{\Phi}$ on $T^{\phi}$. To this end we prove that $\mathrm{Per}_{\mathrm{NP}}(\phi)$ is finite (see Lemma 5.12). We need the following lemmas regarding the intersection of characteristic sets of isometries in $T^{\phi}$.
@@
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$.
Let $v\in VT^{\phi}$ be non-elementary and let $e_{1}$ and $e_{2}$ be distinct edges adjacent to $v$. Let $\phi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. Suppose that there is a representative $\Phi\in\phi$ such that, for every $i=1,2$, we have $\Phi(G_{e_{i}})=G_{e_{i}}$. Then $v$ is rigid and $G_{v}$ is fixed elementwise by some power of $\Phi$.
@@
We first prove that $v$ is rigid. Indeed, suppose towards a contradiction that $v$ is QH with finite fibre. Then $G_{e_{1}}$ and $G_{e_{2}}$ are virtually cyclic. Thus, for every $i\in\{1,2\}$, the automorphism $\Phi$ has a power $\Phi^{k}$ fixing an infinite order element $g_{i}\in G_{e_{i}}$. By Lemma˜5.3 and as the groups $G_{e_{1}}$ and $G_{e_{2}}$ are virtually cyclic, the group $\langle g_{1},g_{2}\rangle$ is a non-elementary subgroup. As $\langle g_{1},g_{2}\rangle\subseteq\mathrm{Per}(\Phi)\cap G_{v}$, this contradicts Lemma˜5.6. Thus, the vertex $v$ is rigid.
By Lemma˜5.4, after taking a power $\Phi^{\ell}$, it acts on $G_{v}$ as global conjugation by an element $g\in G_{v}$.
We claim that, after taking a further power of $\Phi$, the element $g$ is trivial. Indeed, note that, by Lemma˜5.3 the stabiliser of an edge adjacent to $v$ is almost malnormal in $G_{v}$: for every edge $e^{\prime}$ of $T^{\phi}$ adjacent to $v$ and every $g^{\prime}\in G_{v}$, if $g^{\prime}G_{e}g^{\prime-1}\cap G_{e}$ is infinite then $g^{\prime}\in G_{e}$. Moreover, if $e$ and $e^{\prime}$ are two distinct edges adjacent to $v$, then $G_{e}\cap G_{e^{\prime}}$ is finite. Since $\Phi$ preserves $G_{e_{1}}$ and $G_{e_{2}}$, the power $\Phi^{\ell}$ must act by conjugating by an element of the finite intersection $G_{e_{1}}\cap G_{e_{2}}$. This becomes trivial on taking a further power of $\Phi$, which concludes the proof of the claim.
@@
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$. Let $H]=[\mathrm{Per}(\Phi)]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ and let $T_{H}^{\phi}$ be the minimal $H$-invariant subtree of $T^{\phi}$.
@@
Since $T_{H}^{\phi}$ contains an edge, it follows that $H$ is not elliptic in $T^{\phi}$. Thus, $T^{\phi}_{H}$ is the union of the axes of elements of $H$. Then, $e$ is contained in the axis of an element $g$ of $H$. Recalling the bipartite structure of $T^{\phi}$, let $v$ be the rigid or QH vertex adjacent to $e$. Then there exists an edge $e^{\prime}\neq e$ adjacent to $v$ and contained in the axis of $g$. Let $\Phi\in\phi$ be such that $\mathrm{Per}(\Phi)=H$ and let $N\in\mathbb{N}^{*}$ be such that $\Phi^{N}(g)=g$. Since $\Phi^{N}$ is elliptic in $T^{\phi}$ by Remark˜5.7, it fixes pointwise the axis of $g$. In particular, it fixes $e$ and $e^{\prime}$. By Lemma 5.8, we see that $v$ is rigid and that $G_{v}$ is fixed elementwise by a power of $\Phi$. ∎
@@
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$. Let $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ and let $T_{H}^{\phi}$ be the minimal $H$-invariant subtree of $T^{\phi}$.
@@
Suppose first that $T_{H}^{\phi}$ is reduced to a point $v$. Then $H\subseteq G_{v}$ and $v$ is not QH with fibre by Lemma˜5.6. Suppose now that $T_{H}^{\phi}$ is not reduced to a point. Then any vertex $v$ of $T_{H}^{\phi}$ is adjacent to an edge and the result follows from Corollary˜5.9. ∎
@@
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$.
Let $n\in\mathbb{N}^{*}$ and let $\Psi,\Theta\in\phi^{n}$ be such that $F_{\Psi}$ and $F_{\Theta}$ are elliptic isometries of $T^{\phi}$. Suppose that there exist $g,h\in G$ loxodromic in $T^{\phi}$, such that $\Psi(g)=g$, $\Theta(h)=h$ and $\mathrm{Ax}(g)\cap\mathrm{Ax}(h)$ contains an edge $e$.
@@
Therefore, we can apply Lemma 5.8: there exists $N\in\mathbb{N}^{*}$ such that both $\Psi^{N}$ and $\Theta^{N}$ act as the identity on the nonelementary subgroup $G_{v}$. By Lemma˜5.1, up to taking powers of $\Psi$ and $\Theta$, we have $\Psi^{N}=\Theta^{N}$. ∎
@@
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$. The set $\bigcup_{n\in\mathbb{N}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ is finite. Consequently, there exists $N\in\mathbb{N}^{*}$ such that any subgroup whose conjugacy class is in $\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ belongs to a subgroup whose conjugacy class is in $\mathrm{Per}_{\mathrm{NP}}(\phi^{N})$.
@@
Counting elliptic subgroups: Let $H]\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$. Suppose first that every element of $H$ is elliptic in $T^{\phi}$. By Lemma˜4.3, $H$ is elliptic in $T^{\phi}$. Since $H]\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$, the group $H$ is not contained in the stabiliser of a vertex which is elementary. By Corollary˜5.9, it is also not contained in the stabiliser of a QH with finite fibre vertex. Therefore, the group $H$ is contained in the stabiliser of a vertex $v$ which is rigid. By Lemma˜5.6, we have in fact $H=G_{v}$.